yeetantor:

If 1 yard = 3 feet and 1 mile = 5,280 feet, how many yards are there in 2 miles?

Shadow:

\[\frac{ 2 miles }{ 1 } \times \frac{ 5,280feet }{ 1mile } \times \frac{ 1 yard }{ 3feet } = ?\]

Shadow:

Multiply across, cancel like terms. Feet and miles should cancel out, and you should be left with yards.

ramen:

Dimensional Analysis is a great tool.

Shadow:

The way you set this up is to write your unit, then set up the conversion ratios to ensure that it cancels out with the next multiplication. So 2mil/1 has miles on top. So in the next fraction that relates miles to feet, we need to have miles on the bottom and feet on the top, so that when we multiply across, miles is being cancelled as it becomes miles/miles which equals 1. Same concept happens with feet, and then you're left with yards.

yeetantor:

i think i know the answer

yeetantor:

i believe its 3,520 yards

Shadow:

I got the same

yeetantor:

ok then thats the right answer

Shadow:

Yes, good job

yeetantor:

thanks